suppose the price of Good is 1 rupee and price elasticity of demand is 0.2, how much will 10% reduction in quantity demanded increase the price? what will be the new price?


1
Expert's answer
2021-01-12T13:05:14-0500

Price Elasticity of Demand (Ed) = Percentage Change in Quantity Demanded / Percentage Change in Price

(-)0.2 = -10 / Percentage Change in Price

Percentage Change in Price = -10 / -0.2 = 50%

Calculation of New Price: price has increased by 50%

New Price = 1 + 50% of 1 = 1 + 0.50 = 1.50 rupee